# RAVEN: Realtime Accessibility in Virtual ENvironments for Blind and Low-Vision People

### v1

@@ 6.4. Conversational Programming for Personalized Accessibility
Our findings showed marked differences in how categories were valued depending on participants’ visual ability and history of vision. Low-vision participants benefited from color and brightness adjustments, while blind participants often dismissed these as cosmetic or irrelevant. Conversely, blind participants emphasized precise spatial orientation, whereas some low-vision participants preferred richer visual cues. By accommodating both, RAVEN responds to the diverse and sometimes conflicting preferences of BLV users (Szpiro et al., 2016), demonstrating how conversational programming can enable personalized accessibility.
@@ 6.5. Limitations and Future Work
Our study represents an early exploration of conversational programming for accessibility in 3D environments. Several limitations qualify our findings and point to opportunities for future work. First, the system exhibited a non-trivial error rate, meaning it is not yet deployable in real-world applications. Future work should prioritize error prevention and robust verification methods to increase trust and reliability. Second, the system lacked an understanding of object affordances (e.g., distinguishing the sitting surface of a bench from its coordinates), limiting functional interactions. Richer semantic modeling of object properties will be needed to support more realistic accessibility modifications. Third, our evaluation was conducted in a controlled, scenario-based setting with short-term tasks. Longer-term deployments in commercial games or fast-paced contexts (e.g., combat scenarios) are necessary to assess real-world viability.

### v2

@@ 7.5. Conversational Programming for Personalized Accessibility
Our findings showed marked differences in how categories were valued depending on participants’ visual ability and history of vision. Low-vision participants benefited from color and brightness adjustments, while blind participants often dismissed these as cosmetic or irrelevant. Conversely, blind participants emphasized precise spatial orientation, whereas some low-vision participants preferred richer visual cues. By accommodating both, RAVEN responds to the diverse and sometimes conflicting preferences of BLV users (Szpiro et al., 2016), demonstrating how conversational programming can enable personalized accessibility. Rather than treating “blind” and “low-vision” users as two homogeneous groups, our analysis reflects this continuum of abilities and preferences, which further motivates an adaptive, user-directed approach over static presets.
@@ 7.6. Limitations and Future Work
Our study represents an early exploration of conversational programming for accessibility in 3D environments. Several limitations qualify our findings and point to opportunities for future work. First, the system exhibited a non-trivial error rate, meaning it is not yet deployable in real-world applications revTwoand may have negatively impacted user trust and usability during evaluation. Future work should prioritize error prevention and robust verification methods to increase trust and reliability. Second, the system lacked an understanding of object affordances (e.g., distinguishing the sitting surface of a bench from its coordinates), limiting functional interactions. Richer semantic modeling of object properties will be needed to support more realistic accessibility modifications. Third, our evaluation was conducted in a controlled, scenario-based setting with short-term tasks. Although this is comparable to other accessibility evaluations, our sample size (eight BLV participants) limits the statistical power of between-group comparisons, which means our findings should be interpreted primarily as qualitative and exploratory. Finally, longer-term deployments in commercial games or fast-paced contexts (e.g., combat scenarios) are necessary to assess real-world viability.